# PCDreamer: Point Cloud Completion Through Multi-view Diffusion Priors

### v1

@@ Abstract
This paper presents PCDreamer, a novel method for point cloud completion. Traditional methods typically extract features from partial point clouds to predict missing regions, but the large solution space often leads to unsatisfactory results. More recent approaches have started to use images as extra guidance, effectively improving performance, but obtaining paired data of images and partial point clouds is challenging in practice. To overcome these limitations, we harness the relatively view-consistent multi-view diffusion priors within large models, to generate novel views of the desired shape. The resulting image set encodes both global and local shape cues, which is especially beneficial for shape completion. To fully exploit the priors, we have designed a shape fusion module for producing an initial complete shape from multi-modality input (i.e., images and point clouds), and a follow-up shape consolidation module to obtain the final complete shape by discarding unreliable points introduced by the inconsistency from diffusion priors. Extensive experimental results demonstrate our superior performance, especially in recovering fine details.
@@ 1 Introduction
To harness the diffusion priors, we propose PCDreamer, a novel algorithm composed of three key components for high-quality point cloud completion. We, for the first time, introduce diffusion priors in this task. Specifically, in the first module, given the single-view partial point cloud, we utilize a large multi-view diffusion model to ‘dream’ a set of realistic images for free. These multi-view images depict the desired shape, serving as the fuel for the completion. The second module employs the attention mechanism to effectively fuse both the images and partial point cloud observation, leading to an initial full shape that preserves both the global structure and local features. Additionally, to address the inconsistency introduced by the multi-view images, the final module incorporates a confidence-guided shape consolidator, which produces the final complete shape with high fidelity. Extensive evaluation and experiments demonstrate the superior performance of our approach.
@@ 2 Related Work
Recent attention-based approaches, particularly those using vector attention mechanisms [53], have further advanced the field. SnowflakeNet [43] and CRA-PCN [30] introduced novel transformer-based components for improved detail generation. SeedFormer [55] and AnchorFormer [6] both predict shape representations that preserve local information and integrate these representations into the generation process to restore high-fidelity details. Methods like PoinTr [48] use the partial point cloud as a condition, leveraging Transformer architectures to predict missing points. Despite the improvements, Transformer-based approaches still struggle with severely incomplete point clouds due to limited supplementary information.
@@ 3.2 Multi-modality Shape Fusion
Partial Point Cloud Encoder. Due to the nature of single-view scanning, the partial observation is a golden standard for the desired shape, containing global and local geometric priors. We thus adopted a patch-based Transformer encoder to effectively extract geometric features, as shown in Fig. 3 (a). Formally, we first select seed points from the partial point cloud to divide it into patches , based on a fixed radius . Each patch goes into a DGCNN [38] encoder to extract patch-level features, while the center point (i.e., the seed point) of each patch is transformed into a positional encoding via the sinusoidal function. Finally, the Transformer encoder takes as input the concatenated features and outputs a unified feature of the partial point cloud after average pooling, defined as:
Multi-view Image Encoder. Given the posed multi-view depth images, a natural way is to fuse them into 3D and extract features from the fused 3D shape directly. We have proved in the ablation study in Sec. 4.3 that the resulting shape is irregular and noisy, because of the inherent inconsistency of the multi-view generation. To accurately extract global and local shape cues from the images, particularly for the missing regions, while mitigating the impact of inconsistencies, we have designed a patch-based image encoder, as shown in Fig. 3 (b). Specifically, each image goes through a pre-trained ResNet backbone to obtain the image feature, while its corresponding camera pose is further encoded via a two-layer MLP to create the positional encoding. Similarly, the Transformer takes as input the concatenated feature and produces the final multi-view image fusion feature after average pooling, defined as:
@@ 3.3 Shape Consolidation
Our whole network is trained end-to-end with the Chamfer distance loss. Specifically, we measure the discrepancy between the ground truth point cloud and the coarse () and final () point clouds, respectively. Inspired by [47], we exploit the extra hyperbolic Chamfer distance (HyperCD) to combat noise and outliers, defined as:
@@ 4.1 Dataset
PCN dataset. The PCN dataset [49] is a subset of the ShapeNet dataset [5], containing shapes from eight categories. In this study, we constructed our dataset based on PCN. Specifically, given the mesh data of a complete object, we selected a fixed viewpoint (the same as the Wonder3D viewpoint) along the positive z-axis on its unit sphere to obtain a single-view depth image and generated a partial point cloud with 2,048 points through a back-projection method. Additionally, multi-view images were produced from the single-view depth image using the multi-view image generation module (Sec. 3.1). We uniformly sampled 16,384 points from the mesh surface to obtain the corresponding ground truth point cloud for each object. In total, we collected 30974 samples and split the dataset into training, validation, and test sets, the ratio is the same as [49].
ShapeNet-55 dataset. Unlike the PCN dataset [49], ShapeNet-55 [48] includes shapes from 55 categories. We apply the same construction method as aforementioned to create a dataset.
Evaluation metrics. Following recent works [55, 56], we adopt Chamfer Distance (CD), Density-aware CD (DCD) [42], and F1-Score as evaluation metrics, and report the version of CD for PCN dataset and the version of CD for Shapenet-55 dataset.
@@ 4.3 Ablation study
Is confidence filtering helpful? To evaluate the impact of our proposed confidence-based filtering operation, we conducted an ablation study by removing it from our method. As shown in Fig. 8, the red box in (b) highlights noisy points that are assigned low confidence scores by our method. By applying the filtering strategy, as shown in (c), these noisy points are effectively removed, resulting in a consolidated and improved completion result. Quantitative results in Tab. 5 further demonstrate the effectiveness of the confidence filtering operation.
@@
4.3.1 Discussion
@@
Limitation. Multi-view depth images can guide point cloud completion, but there are limitations in generating these images. As shown in the Fig. 9, the slender structure of the earphone exhibits significant inconsistencies across different generated views, resulting in poor point cloud completion when used as guidance. In cases like this, where the information visible in a single view is extremely limited, the completion results are often unsatisfactory.
@@ 5 Conclusion
In this paper, we have presented a novel approach for point cloud completion. We, for the first time, introduce diffusion priors in this task to provide imagined full-shape cues for free. From this proof of concept, we believe that diffusion priors can be useful for other point cloud processing tasks, and we hope PCDreamer can inspire future research in this direction.
@@
In this supplementary, we present more visual results on the PCN [49] and ShapeNet-55 [48] datasets. We also present additional visual examples, comparing our approach against more single-view 3D generation methods. Besides, to further demonstrate the generalization ability of our method in real-world scans, we also conduct experiments on the ScanNet dataset [7], which contains partial point clouds extracted from RGB-D scans.
Additional results on Real-world Scans.
We evaluated our method on real-world scanned data from the ScanNet dataset [7] to assess its performance. Specifically, we selected depth images of individual objects and their corresponding partial point clouds from ScanNet. Using our multi-view generation module, we obtained multi-view depth images. These were then input into our model, which was pre-trained on the PCN dataset, to produce the completion results. As shown in Fig. 1, our method yields reasonable and high-quality results on real scanned data, demonstrating its effectiveness and robustness in handling real-world inputs.
Visual comparison.
In Fig. 2, we present additional comparisons with other 3D model generation methods [39, 34]. Moreover, we provide further examples from the PCN dataset for comprehensive evaluation. As illustrated in the figure, the lack of extra views, the randomness in the generation process, and the inherent inconsistencies in multi-view images often prevent single-view 3D model generation methods from effectively capturing geometric details. Consequently, these models may produce shapes that, while plausible, exhibit significant deviations from the ground truth, such as variations in the width of the sofa. Therefore, multi-view images are more appropriate as auxiliary cues for point cloud completion, although directly utilizing them for point cloud completion typically yields suboptimal results.
@@
Figs. 3 and 4, each presents an extra six results from the PCN and the ShapeNet-55 datasets, respectively. For both datasets, the first three cases in the first three rows possess high-quality and consistent multi-view depth images, whereas depth images with lower quality and less consistency are generated for the last three examples. Specifically, the high randomness of the generation process, combined with limitations of the initial view, can result in multi-view depth images exhibiting inconsistencies across views (e.g., the lid of the Dustbin), missing geometric details(e.g., the Helmet), and the presence of noisy regions (e.g., the leg of the Table). Furthermore, shapes imagined by the generative model may exhibit scale discrepancies relative to their real-world counterparts (e.g., the Sofa and the Table). In addition, the depth estimation process may also occasionally produce suboptimal outcomes (e.g., the Dishwasher). However, the proposed confidence-based shape consolidator effectively addresses these issues by eliminating unreliable points caused by inconsistencies in diffusion priors. As a result, our models produce accurate and reasonable outputs, as demonstrated in the second and fourth columns of the last three examples.

### v2

@@ Abstract
This paper presents PCDreamer, a novel method for point cloud completion. Traditional methods typically extract features from partial point clouds to predict missing regions, but the large solution space often leads to unsatisfactory results. More recent approaches have started to use images as extra guidance, effectively improving performance, but obtaining paired data of images and partial point clouds is challenging in practice. To overcome these limitations, we harness the relatively view-consistent multi-view diffusion priors within large models, to generate novel views of the desired shape. The resulting image set encodes both global and local shape cues, which are especially beneficial for shape completion. To fully exploit the priors, we have designed a shape fusion module for producing an initial complete shape from multi-modality input (i.e., images and point clouds), and a follow-up shape consolidation module to obtain the final complete shape by discarding unreliable points introduced by the inconsistency from diffusion priors. Extensive experimental results demonstrate our superior performance, especially in recovering fine details.
@@ 1 Introduction
To harness the diffusion priors, we propose PCDreamer, a novel algorithm composed of three key components for high-quality point cloud completion. Specifically, in the first module, given the single-view partial point cloud, we utilize a large multi-view diffusion model to ‘dream’ a set of realistic images for free. These multi-view images depict the desired shape, serving as the fuel for the completion. The second module employs the attention mechanism to effectively fuse both the images and partial point cloud observation, leading to an initial full shape that preserves both the global structure and local features. Additionally, to address the inconsistency introduced by the multi-view images, the final module incorporates a confidence-guided shape consolidator, which produces the final complete shape with high fidelity. Extensive evaluation and experiments demonstrate the superior performance of our approach.
@@ 2 Related Work
Recent attention-based approaches, particularly those using vector attention mechanisms [55], have further advanced the field. SnowflakeNet [45] and CRA-PCN [31] introduced novel transformer-based components for improved detail generation. SeedFormer [57] and AnchorFormer [6] both predict shape representations that preserve local information and integrate these representations into the generation process to restore high-fidelity details. Methods like PoinTr [50] use the partial point cloud as a condition, leveraging Transformer architectures to predict missing points. Despite the improvements, Transformer-based approaches still struggle with severely incomplete point clouds due to limited supplementary information. SDS-Complete [17] employs semantic priors for point cloud completion, but it suffers from low efficiency and noisy outputs. We compare our method with it qualitatively in the supplementary.
@@ 3.2 Multi-modality Shape Fusion
Partial Point Cloud Encoder. Due to the nature of single-view scanning, the partial observation is a golden standard for the desired shape, containing global and local geometric priors. We thus adopted a patch-based Transformer encoder to effectively extract geometric features, as shown in Fig. 3 (a). Formally, we first select seed points from the partial point cloud to divide it into patches , based on a fixed radius . Each patch goes into a DGCNN [40] encoder to extract patch-level features, while the center point (i.e., the seed point) of each patch is transformed into a positional encoding via the sinusoidal function (denoted as Sins). Finally, the Transformer encoder takes as input the concatenated features and outputs a unified feature of the partial point cloud after average pooling, defined as:
Multi-view Image Encoder. Given the posed multi-view depth images, a natural way is to fuse them into 3D and extract features from the fused 3D shape directly. We have proved in the ablation study in Sec. 4.3 that the resulting shape is irregular and noisy, because of the inherent inconsistency of the multi-view generation. To accurately extract global and local shape cues from the images, particularly for the missing regions, while mitigating the impact of inconsistencies, we have designed a patch-based image encoder, as shown in Fig. 3 (b). Specifically, each image goes through a ResNet backbone to obtain the image feature, while its corresponding camera pose is further encoded via a two-layer MLP to create the positional encoding. Similarly, the Transformer takes as input the concatenated feature and produces the final multi-view image fusion feature after average pooling, defined as:
@@ 3.3 Shape Consolidation
Our network is trained end-to-end with the Chamfer distance loss. Specifically, we measure the discrepancy between the ground truth point cloud and the coarse () and final () point clouds, respectively. Inspired by [49], we exploit the extra hyperbolic Chamfer distance (HyperCD) to combat noise and outliers, defined as:
@@ 4.1 Dataset
PCN dataset. The PCN dataset [51] is a subset of the ShapeNet dataset [5], containing shapes from eight categories. In this study, we constructed our dataset based on PCN. Specifically, given the mesh data of a complete object, we selected a fixed viewpoint along the positive z-axis on its unit sphere to obtain a single-view depth image and generated a partial point cloud with 2,048 points through a back-projection method. The adoption of a fixed viewpoint is mainly because the large diffusion models (e.g., SVD-MV and Wonder3D) require such an input viewpoint configuration. Additionally, multi-view images were produced from the single-view depth image using the multi-view image generation module (Sec. 3.1). We uniformly sampled 16,384 points from the mesh surface to obtain the corresponding ground truth point cloud for each object. In total, we collected 30974 samples and split the dataset into training, validation, and test sets, the ratio is the same as [51].
@@ 4.2 Comparisons
| Methods | 34 seen categories | 21 unseen categories | ||||
| CD | DCD | F1 | CD | DCD | F1 | |
| SVDFormer | 0.839 | 0.625 | 0.402 | 1.129 | 0.617 | 0.348 |
| CRA-PCN | 0.818 | 0.591 | 0.387 | 1.277 | 0.601 | 0.317 |
| Ours(SVD-MV) | 0.765 | 0.562 | 0.396 | 1.040 | 0.574 | 0.339 |
| Ours(Wonder3D) | 0.758 | 0.573 | 0.403 | 1.094 | 0.582 | 0.352 |
Results on the ShapeNet-34 Dataset. To further evaluate the performance of our method on unseen object categories, we also conduct experiments on the ShapeNet-34 dataset [50]. Tab. 3 displays the statistical results, where PCDreamer maintains high completion quality on unseen categories (see the CD and DCD metrics) compared against state-of-the-art methods [58, 31], demonstrating moderately strong generalization ability. Note that we process the data as stated in Sec. 4.1.
@@ 4.3 Ablation study
Is confidence filtering helpful? To evaluate the impact of our proposed confidence-based filtering operation, we conducted an ablation study by removing it from our method. As shown in Fig. 8, the red box in (b) highlights noisy points that are assigned low confidence scores by our method. By applying the filtering strategy, as shown in (c), these noisy points are effectively eliminated. The remaining points with high confidence are fused with the input and are further consolidated to obtain an improved completion result (Fig. 8(d)). Quantitative results in Tab. 6 further demonstrate the effectiveness of the confidence filtering operation.
4.4 Discussions
In the following, we primarily discuss the impact of the number of views, the hyperparameter configuration of filtering, as well as the limitation, while leaving more discussions on handling real-world LiDAR scans and general partial point clouds in the supplementary.
@@ 4.4 Discussions
Limitation. Multi-view depth images can guide point cloud completion, but there are limitations in generating these images. As shown in Fig. 9, the slender structure of the earphone exhibits significant inconsistencies across different generated views, resulting in poor point cloud completion when used as guidance. In cases like this, where the information visible in a single view is extremely limited, the completion results are often unsatisfactory.
@@ 5 Conclusion
In this paper, we have presented a novel approach for point cloud completion. We introduce diffusion priors in this task to provide imagined full-shape cues for free. From this proof of concept, we believe that diffusion priors can be useful for other point cloud processing tasks, and we hope PCDreamer can inspire future research in this direction.
@@
In this supplementary, we present experiments on additional datasets (i.e., ScanObjectNN [37], KITTI [8], ShapeNet-55 [50]) to demonstrate the generalization ability of our method on real-world scans, as well as randomly cropped general partial points clouds. Besides, we show complementary visual results on PCN [51] and ShapeNet-55 [50] datasets and present additional visual examples when comparing our approach against SDS-Complete [17] and other single-view 3D generation methods.
Real-world scans and general partial point clouds.
As stated in Sec. 1 in the paper, completing a single-view partial point cloud is actively studied and much more challenging than completing a general one, since it usually misses more than half of the points (see the chair and lamp in Fig. 1 in the main paper). Due to the unique setting, we thus design our PCDreamer, dedicated to completing a single-view partial point cloud faithfully. However, thanks to the robustness of the large diffusion models and our fuser and consolidator, PCDreamer can work on general partial clouds well. Following the convention, we did not emphasize and report this advantage in the paper, but preliminary visual results are presented in Fig. A1, where PCDreamer successfully completes both LiDAR point clouds and randomly cropped ShapeNet55 point clouds. Indeed, a perfect single-view depth image is not feasible in this case, we thus select the most informative view and obtain the ‘incomplete’ depth image serving as the input of our method. To mitigate the incompleteness and potential quality degradation in the initial depth map, we employ a larger guidance scale (e.g., 9.0) along with more complex text prompts in ControlNet. This strategy effectively enhances the quality of the generated RGB images. Extending our method to handle arbitrary multi-view depth images is a promising direction, we leave it for future work.
Comparison with SDS-Complete [17].
We have conducted a visual comparison with SDS-Complete, which exploits diffusion priors with SDS optimization. The results in Fig. A2 demonstrate that while SDS-Complete is capable of recovering the overall shape, it fails to capture fine details, such as the leg and the armrest of the chair.
Additional Comparison with 3D generation methods.
In Fig. A3, we present additional comparisons with other 3D model generation methods [41, 35]. Moreover, we provide further examples from the PCN dataset for comprehensive evaluation. As illustrated in the figure, the lack of extra views, the randomness in the generation process, and the inherent inconsistencies in multi-view images often prevent single-view 3D model generation methods from effectively capturing geometric details. Consequently, these models may produce shapes that, while plausible, exhibit significant deviations from the ground truth, such as variations in the width of the sofa. Therefore, multi-view images are more appropriate as auxiliary cues for point cloud completion, although directly utilizing them for point cloud completion typically yields suboptimal results.
@@
Figs. A4 and A5, each presents an extra six results from the PCN and the ShapeNet-55 datasets, respectively. For both datasets, the first three cases in the first three rows possess high-quality and consistent multi-view depth images, whereas depth images with lower quality and less consistency are generated for the last three examples. Specifically, the high randomness of the generation process, combined with limitations of the initial view, can result in multi-view depth images exhibiting inconsistencies across views (e.g., the lid of the Dustbin), missing geometric details(e.g., the Helmet), and the presence of noisy regions (e.g., the leg of the Table). Furthermore, shapes imagined by the generative model may exhibit scale discrepancies relative to their real-world counterparts (e.g., the Sofa and the Table). In addition, the depth estimation process may also occasionally produce suboptimal outcomes (e.g., the Dishwasher). However, the proposed confidence-based shape consolidator effectively addresses these issues by eliminating unreliable points caused by inconsistencies in diffusion priors. As a result, our models produce accurate and reasonable outputs, as demonstrated in the second and fourth columns of the last three examples.